# Quantum machines using $\mathrm{Cu}_{3}$-like compounds modeled by Heisenberg antiferromagnetic in a triangular ring

### v2

@@ VI Conclusions
The MCE impact on quantum machine operation and efficiency is significant. We explored the Carnot, Otto, and Stirling machines within a ${\rm Cu}_{3}$-like system as the working substance, focusing on how external magnetic fields and temperature affect their performance assuming reversible process. The Carnot machine operates as a heat engine for $B_{1}\lesssim B_{0}$ and as a refrigerator for $B_{1}\gtrsim B_{0}$, with a shift in mode near the MCE. The Otto machine is versatile, mainly functioning as a thermal accelerator for $B_{1}<B_{0}$ and as a refrigerator for $B_{1}>B_{0}$, with a small region where it acts as a heat engine and heater. The Stirling machine alternates between a heat engine and thermal accelerator near the MCE, operating primarily as a heat engine for $B_{1}>B_{0}$ and as a refrigerator for $B_{1}<B_{0}$. We also evaluated their thermal efficiencies for each machines. This study enhances the understanding of quantum heat machines, especially in low-temperature regimes where the MCE is crucial, offering insights into practical applications of the ${\rm Cu}_{3}-X$ system in quantum thermodynamics.
@@ Acknowledgements.
O. R. and M. Rojas thanks CNPq and FAPEMIG for partial financial support. M. Rojas acknowledges FAPEMIG Grant APQ-02226-22.

### v3

@@ VI Conclusion
The MCE impact on quantum machine operation and efficiency is significant. We explored the Carnot, Otto, and Stirling machines within a ${\rm Cu}_{3}$-like system as the working substance, focusing on how external magnetic fields and temperature affect their performance assuming reversible process. The Carnot machine operates as a heat engine for $B_{1}\lesssim B_{0}$ and as a refrigerator for $B_{1}\gtrsim B_{0}$, with a shift in mode near the MCE. The Otto machine is versatile, mainly functioning as a thermal accelerator for $B_{1}<B_{0}$ and as a refrigerator for $B_{1}>B_{0}$, with a small region where it acts as a heat engine and heater. The Stirling machine alternates between a heat engine and thermal accelerator near the MCE, operating primarily as a heat engine for $B_{1}>B_{0}$ and as a refrigerator for $B_{1}<B_{0}$. We also evaluated their thermal efficiencies for each machines. This study enhances the understanding of quantum heat machines, especially in low-temperature regimes where the MCE is crucial, offering insights into practical applications of the ${\rm Cu}_{3}-X$ system in quantum thermodynamics. Although our analysis focused solely on the compound ${\rm Cu}_{3}$-As, the second compound ${\rm Cu}_{3}$-Sb will also exhibit similar behavior for Carnot, Otto, and Stirling machines, with no relevant changes worth commenting on, as analyzed in Ref.[52]. We did not include any plots here to avoid repetitive behavior.
Future research could focus on optimizing quantum machines using ${\rm Cu}_{3}$-like systems through material tuning, analyzing non-ideal quantum conditions, and expanding applications in quantum refrigeration, assuming non-reversible quantum machines.
@@ Acknowledgements.
O. R. and M. Rojas thank CNPq and FAPEMIG for partial financial support. M. Rojas acknowledges FAPEMIG Grant APQ-02226-22.